# Classification of connection graphs of global attractors for $S^{1}$-equivariant parabolic equations

### v2

@@ 1 Introduction
where the nonlinearity $f=f(u,u_{x})$ satisfies appropriate smoothness and dissipative conditions. Then, (1.1) generates a semiflow in the Sobolev space $X=H^{s}(S^{1}),s>\frac{3}{2}$, which possesses a nonempty compact global attractor $\mathcal{A}\subset X$. For details see [11] and also [3, 26, 27] for initial references. As general references, see [18, 29] for semiflows, and [16, 4, 17] for global attractors.
@@
Hyperbolicity here concerns the sets of equilibria and nonhomogeneous periodic solutions of the semilinear parabolic PDE (1.1). These are, respectively, the sets $\mathcal{E}$ and $\mathcal{P}$:
(1.4) $\mathcal{E}$ =$=\{{\bf e}_{j}=(e_{j},0):1\leq j\leq n\}\ ,$
@@
where $p=v_{x}$ and $q$ denotes the number of nonhomogeneous periodic solutions properly $x$-shifted so that $v_{j}(0)=\min_{x\in[0,2\pi]}v_{j}(x)$. We point out that these solutions may have very large Morse indices $i({\bf e}_{j})=\mathop{\mathrm{dim}}\nolimits W^{u}({\bf e}_{j})$ and $i({\bf v}_{j})=\mathop{\mathrm{dim}}\nolimits W^{u}({\bf v}_{j})-1$, see [11]. In contrast, the equilibria of the related ODE (1.2) has only alternating saddles and centers, all in one-to-one correspondence with the equilibria of (1.1). For example, the first saddle corresponds to the first stable equilibrium ${\bf e}_{1}$ and the last saddle corresponds to the last stable equilibrium ${\bf e}_{n}$.
@@
Notice that a nonhomogeneous stationary solution $v=v(\cdot)$ is not isolated, since all its shifted copies $v(\cdot+\theta),\theta\in\mathbb{R}$, are also $2\pi$-periodic solutions of (1.2). Hence, it is not hyperbolic in stricto sensu. To overcome this inconvenience, hyperbolicity in this case is understood as normal hyperbolicity, [12, 31].
@@
If all critical elements in $\mathcal{E}\cup\mathcal{P}$ are strictly hyperbolic, i.e. $\mathcal{P}$ does not contain nonhomogeneous stationary solutions, the global attractor $\mathcal{A}$ is a Sturm attractor (see [13]) due to the decay property of the zero number, [27]. In addition it has the Morse-Smale property, [7, 20]. In this case all the heteroclinic orbit connections between equilibria and periodic orbits of $\mathcal{A}$, which all together compose the global attractor, are determined by a Sturm permutation, [13, 31]. See also [14] for the initial results on Sturm permutations. Unfortunately, in the $S^{1}$-equivariant case like the reversible class of nonlinearities $f(u,u_{x})=f(u,-u_{x})$, we are missing a proper Morse-Smale Theorem. Therefore, our results must be restricted to connection equivalence of global attractors postponing orbit equivalence to further research studies. This observation also applies to the statement of Theorem 3 in [31] which needs a correction, addresing connection equivalence of global attractors instead of orbit equivalence.
The existence of two special homotopies of $f(u,u_{x})$, which greatly simplify our task, was established by Fiedler, Rocha and Wolfrum in [12]. The first homotopy freezes all rotating orbits to speed $c=0$, by introducing corotating coordinates with nonuniform wave speed $c=c(v,v_{x})$. This homotopy preserves hyperbolicity of all rotating periodic orbits and their spatial periods, therefore preserves the heteroclinic connectivity between them. In addition, the phase portrait of (1.3) with $c=c(v,v_{x})$ is integrable in the phase plane region corresponding to the cyclicity set $\mathcal{C}$, i.e. the region of all spatially periodic orbits, see [12]. Therefore, for the nonlinearity $f_{0}(u,u_{x})$ with all rotating waves frozen, we obtain a period map $T:\mathcal{D}\subset\mathbb{R}\rightarrow\mathbb{R}_{+}$ where the domain $\mathcal{D}$ corresponds to (the $u$-values) of the cyclicity set $\mathcal{C}$.
The second homotopy $f_{\tau}(u,u_{x}),\tau\in[0,1]$, symmetrizes the phase portrait of (1.2) with respect to the $v$-axis obtaining a reversible nonlinearity $f(u,v)=f(u,-v)$ which is even in $v$. This homotopy also preserves the hyperbolicity of all homogeneous equilibria and frozen rotating waves. Moreover, due to the integrability of (1.2), the period map obtained for $f_{0}(u,u_{x})$ is invariant under this second homotopy.
@@
Our objective is the classification of the connection graphs of all Sturm global attractors of flows generated by (1.1). This is achieved using the lap signature class introduced in [31, 12]. The lap signature of a period map $T=T(u_{0})$ consists of the set of period lap numbers of the $2\pi$-periodic solutions of (1.2) endowed with a total order derived from the nesting of the periodic orbits in the phase space $(v,v_{x})$, and their regular parenthesis structure (called regular bracket structure in [23, 31]). We notice that the period lap number $\ell(v)$ of $v\in\mathcal{P}$ is half of the zero number: $z(v)=2\ell(v)$, see [12].
Let $K=(n-1)/2+l$ denote the number of connected regions of the cyclicity set $\mathcal{C}$. Here $n$ denotes the odd number of equilibria and $l$ the number of annular regions. Notice that each equilibrium of (1.1) corresponds to a center or a saddle point of the phase portrait of (1.2). The centers and saddles alternate for increasing values of $u_{0}$, starting and ending with saddles. Then $(n-1)/2$ is the number of punctured disks encircling the centers. Moreover, $l$ is the number of annular regions which surround more than one single center. This implies that $20\leq l\leq(n-3)/$ 2. See [12].
Finally, a reversible nonlinearity $f=f(u,u_{x})$ is of simple type if $l=0$, i.e. if there are no annular regions. This is a slightly more restrictive definition then the one used in [31]. If $l=0$ then each $2\pi$-periodic orbit of (1.2) encircles exactly one center in the phase plane $(v,v_{x})$.
@@
If $f=f(u,u_{x})\in\mathcal{R}$ and $g=g(u,u_{x})\in\mathcal{R}$ belong to the same lap signature class (see [31, 12]), then there exists a homotopy in $\mathcal{R}$ between $f$ and $g$ which preserves hyperbolicity (H) of all the homogeneous equilibria and (normally hyperbolic) frozen rotating waves.
@@
The construction of the homotopy is very delicate and needs some clarification in what concerns the proof in [31], which is restricted to nonlinearities of simple type. In the next Sect. 2 we change some aspects of the proof of Theorem 3 in [31] making it simpler and amenable to a generalization to nonlinearities of non-simple type.

### v5

@@ 1 Introduction
where the nonlinearity $f:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is $C^{2}$ and dissipative. Sufficient dissipative conditions are boundedness, $|f(\cdot,\cdot)|\leq C_{0}$, and a sign condition $uf(u,0)<0$ for all large $|u|>K_{0}$. For less restrictive conditions see [1, 31]. Then, (1.1) generates a semiflow in the Sobolev space $X=H^{s}(S^{1}),s>\frac{3}{2}$, which possesses a nonempty compact global attractor $\mathcal{A}\subset X$. For details see [14] and also [4, 30, 31] for initial references. As general references, see [21, 33] for semiflows, and [19, 5, 20] for global attractors.
@@
Hyperbolicity here concerns the sets of homogeneous equilibria and nonhomogeneous time periodic solutions (rotating waves) of the semilinear parabolic PDE (1.1). These are, respectively, the sets $\mathcal{E}$ and $\mathcal{P}$ which are finite due to hyperbolicty (H) and the compacteness of $\mathcal{A}$, [13, 14]:
(1.4) $\mathcal{E}$ =$=\{e_{j}:f(e_{j},0)=0,1\leq j\leq n\}\ ,$
@@
where $p=v_{x}$, $v_{j}$ denotes the rotating wave profiles, i.e. solutions of (1.3), and $q$ denotes the number of nonhomogeneous periodic solutions properly $x$-shifted so that $v_{j}(0)=\min_{x\in S^{1}}v_{j}(x)$. For simplicity we include in $\mathcal{P}$ also the nonhomogeneous stationary waves. We point out that the critical solutions may have very large Morse indices $i(e_{j})=\mathop{\mathrm{dim}}\nolimits W^{u}(e_{j})$ and $i(v_{j})=\mathop{\mathrm{dim}}\nolimits W^{u}(v_{j})-1$, see [14]. Here $W^{u}(\cdot)$ denotes the unstable manifold of an equilibrium or periodic orbit, [23, 20, 21]. In contrast, the equilibria of the related ODE (1.2), ${\bf e_{j}}=(e_{j},0)$ and ${\bf v_{j}}=(v_{j},p_{j})$, has only alternating saddles and centers, all in one-to-one correspondence with the equilibria of (1.1). For example, the first saddle corresponds to the first stable equilibrium ${\bf e}_{1}$ and the last saddle corresponds to the last stable equilibrium ${\bf e}_{n}$.
@@
Notice that a nonhomogeneous stationary solution $v=v(\cdot)$ is not isolated, since all its shifted copies $v(\cdot+\vartheta),\vartheta\in S^{1}$, are also $2\pi$-periodic solutions of (1.2). Hence, a nonhomogeneous stationary solution is not hyperbolic in stricto sensu. To overcome this inconvenience, hyperbolicity here is understood as normal hyperbolicity, [15, 35], much like in the case of a periodic orbit.
@@
The global attractor $\mathcal{A}$ is called a Sturm attractor (see [12]) due to the decay property (2.5) of the zero number, [31]. If all the critical elements in $\mathcal{E}\cup\mathcal{P}$ are strictly hyperbolic, i.e. $\mathcal{P}$ does not contain nonhomogeneous stationary solutions, the Sturm attractor has the Morse-Smale property, [10, 24]. In this case all the heteroclinic orbit connections between equilibria and periodic orbits of $\mathcal{A}$, which all together compose the global attractor, are determined by a Sturm permutation, [16, 35]. See also [17] for the initial results on Sturm permutations. Unfortunately, for the reversible class of nonlinearities $f(u,v)=f(u,-v)$ included in the present $S^{1}$-equivariant case, we are missing a proper Morse-Smale Theorem. Therefore, our results must be restricted to connection equivalence of global attractors postponing orbit equivalence to further research studies. This observation also applies to the statement of Theorem 3 in [35] which should address connection equivalence of global attractors instead of orbit equivalence.
The phase portrait of (1.2) for a reversible nonlinearity $f(u,v)=f(u,-v)$ (even in $v$) has all rotating waves frozen, i.e. they are nonhomogeneous equilibria. In addition, (1.2) is integrable in the phase plane region corresponding to the cyclicity set $\mathcal{C}$, i.e. the region of all spatially periodic orbits, see [15]. Hence, for the reversible nonlinearity $f(u,u_{x})$ we obtain a period map $T:\mathcal{D}\subset\mathbb{R}\rightarrow\mathbb{R}_{+}$ where the domain $\mathcal{D}$ corresponds to (the $u$-values) of the cyclicity set $\mathcal{C}$.
The period map $T=T(u_{0})$, also called time map, is essential for the characterization of all $2\pi$-periodic solutions of autonomous planar Hamiltonian equations, see [15, 34]. In our setting, the period map is used for the characterization of all the equilibria of (1.1), in particular the frozen rotating waves.
@@
Our objective is the classification of the connection graphs of all Sturm global attractors of flows generated by (1.1). This is achieved using the lap signature class introduced in [35, 15]. The lap signature of a period map $T=T(u_{0})$ consists of the set of period lap numbers of the $2\pi$-periodic solutions of (1.2) endowed with a total order derived from the nesting of the periodic orbits in the phase space $(v,v_{x})$, and their regular parenthesis structure (called regular bracket structure in [27, 35]). We notice that the period lap number $\ell(v)$ of $v\in\mathcal{P}$ is half of the zero number: $z(v)=2\ell(v)$, see [15]. The lap signature class of a period map $T=T(u_{0})$ is the set of period maps with the same lap signature of $T=T(u_{0})$.
@@
Let $f=f(u,u_{x})\in\mathcal{R}$ and $g=g(u,u_{x})\in\mathcal{R}$ denote two reversible nonlinearities with period maps in the same lap signature class. Then the corresponding global attractors are connection equivalent,
(1.5) $\mathcal{A}_{f}\sim\mathcal{A}_{g}\ .$
Here (1.5) means connection equivalence between global attractors, [12, 13, 35]. This result is not immediate because all the critical elements in $\mathcal{P}$ fail the hyperbolicity condition (H) which would entail the automatic transversality of stable and unstable manifolds and the strong Morse-Smale property. Instead, we will address the automatic transversality of center stable and center unstable manifolds in the next Sect. 2. Then, Theorem 1 will follow from:
Theorem 2.
@@
If $f=f(u,u_{x})\in\mathcal{R}$ and $g=g(u,u_{x})\in\mathcal{R}$ belong to the same lap signature class (see [35, 15]), then there exists a global collective homotopy in $\mathcal{R}$ between $f$ and $g$ which preserves hyperbolicity (H) of all the homogeneous equilibria and (normally hyperbolic) frozen rotating waves.
@@
The construction of this global collective homotopy is very delicate and needs some clarification in what concerns the proof in [35], which is restricted to nonlinearities of simple type. In Sect. 3 we change and refine some aspects of the proof of Theorem 3 in [35] making it simpler and amenable to a generalization to nonlinearities of non-simple type.